# Manuel Faysse1,5  Patrick Fernandes6,8,11  Nuno M. Guerreiro2,5,6,8 António Loison1  Duarte M. Alves6,8  Caio Corro9  Nicolas Boizard4,5 João Alves2  Ricardo Rei2,7,8  Pedro H. Martins2  Antoni Bigata Casademunt10

### v4

@@ 5.4 Model limitations
Multilingual performance. CroissantLLM is mostly a French and English model. Code performance is relatively limited, and although some amount of data from other languages is included within the SlimPajama training set, out-of-the-box performance in other languages is not to be expected, although some European languages do work quite well.
Hallucinations. CroissantLLM can hallucinate (Ji et al., 2023; Guerreiro et al., 2023) and output factually incorrect data, [^60] As an example, prompted with ”Which French club won the UEFA Champions League ?”, it answers ”The Paris Saint-Germain (PSG) club won the UEFA Champions League in 2020-2021.” especially regarding complex topics. This is to be expected given the small model size, and hallucination rates seem inferior to most models of the same size category although no quantitative assessments have been conducted outside of MT-Bench experiments.
@@ D.1 Methodology
For generative tasks, we simply generate with the default settings, which is greedy sampling. We acknowlege CroissantLLM works best with higher temperature values but did not want to introduce stochasticity to the evaluation. We also limit each benchmark task to 2000 samples at most, to shorten evaluation time. All evaluations are reproducible through the code at https://github.com/EleutherAI/lm-evaluation-harness.
@@ D.2 MT-Bench
Turn 1 (Figure 11) and Turn 2 (Figure 12) results are shown. We notice small models, struggle with reasoning based tasks and contrained generation imposed by Turn 2 prompts. Figures 14 and 15 compare our results on small language models to other common bigger models. [^65] Results in French for models with sizes over 7B parameters were extracted from https://huggingface.co/datasets/bofenghuang/mt-bench-french and results in English are from https://huggingface.co/spaces/lmsys/mt-bench/tree/main/data/mt_bench/model_judgment
@@ D.3 Bias Assessment
We assess bias through CROWS (Nangia et al., 2020), the Crowdsourced Stereotype Pairs benchmark that cover stereotypes dealing with nine types of bias, like race, religion, and age and report results in Table 16. We find CroissantLLM is in line, or slightly less biased than other models, notably in French.
@@
Table 16: Code Data mix

### v5

@@ 5.4 Model limitations
Hallucinations. CroissantLLM can hallucinate (Ji et al., 2023; Guerreiro et al., 2023) and output factually incorrect data, [^65] As an example, prompted with ”Which French club won the UEFA Champions League ?”, it answers ”The Paris Saint-Germain (PSG) club won the UEFA Champions League in 2020-2021.” especially regarding complex topics. This is to be expected given the small model size, and hallucination rates seem inferior to most models of the same size category although no quantitative assessments have been conducted outside of MT-Bench experiments. Veering away from generative use cases, CroissantLLM has also been adapted as an embedding model  [^66] https://huggingface.co/manu/sentence_croissant_alpha_v0.4 and recent work has since been conducted in assessing confidence in the retrieval scores associated to such models (Gisserot-Boukhlef et al., 2024).
@@ 6 Foundation Model Transparency Index
We would like to give a warm thanks to Stephane Requena for allowing us access to Jeanzay, Rémi Lacroix, and Etienne Malaboeuf for the debugging and technical support, and the management from CentraleSupelec for supporting this project from the start: Renaud Monnet and Paul-Henry Cournede.
@@ D.1 Methodology
For generative tasks, we simply generate with the default settings, which is greedy sampling. We acknowledge CroissantLLM works best with higher temperature values but did not want to introduce stochasticity to the evaluation. We also limit each benchmark task to 5000 samples at most, to shorten evaluation time. All evaluations are reproducible through the code at https://github.com/EleutherAI/lm-evaluation-harness.
@@ D.2 MT-Bench
Turn 1 (Figure 11) and Turn 2 (Figure 12) results are shown. We notice small models, struggle with reasoning based tasks and contrained generation imposed by Turn 2 prompts. Figures 13 and 14 compare our results on small language models to other common bigger models. [^71] Results in French for models with sizes over 7B parameters were extracted from https://huggingface.co/datasets/bofenghuang/mt-bench-french and results in English are from https://huggingface.co/spaces/lmsys/mt-bench/tree/main/data/mt_bench/model_judgment
@@ D.3 Bias Assessment
We assess bias through CROWS (Nangia et al., 2020; Névéol et al., 2022), the Crowdsourced Stereotype Pairs benchmark that cover stereotypes dealing with nine types of bias, like race, religion, and age and report results in Table 15. We find CroissantLLM is in line, or slightly less biased than other models, notably in French.
@@
Table 15: Bias Evaluation through Crows-Pairs dataset assessed by likelihood difference.
D.4 MMLU Results
[…]
Appendix E Terminology
Large Language Models. According to the definition proposed by Rogers & Luccioni (2024), Large Language Models (LLMs) (1) have the capacity to model and generate text, (2) are pretrained on large amounts of text data (over 1B tokens according to the proposed threshold) (3) enable transfer learning through finetuning or prompting. Under these criteria, CroissantLLM can largely be considered a large language model despite its small size relative to current best-in-class LLMs.
Small Language Models. Small Language Models (SLMs) is a term that has appeared more recently and that refers to LLMs with a relatively low parameter count, often optimized for speed or to enable inference with local or lower-end hardware.
Effective Capacity Ratio. Introduced in Fernandes et al. (2023), the effective parameters of a multilingual model in a particular language is the number of non-embedding parameters a monolingual model would require to match the language performance of the multilingual model. The intuition is that a multilingual model splits its capacity between different languages, so a monolingual model would reach similar performances on a specific language with a fraction of the parameters. We further compute the effective capacity ratio by dividing the effective parameters by the multilingual model size. In our joint scaling laws experiments (Figure 4), we show that the models trained with equal ratios of English to French have an effective capacity ratio of over 82% in the French language, indicating that training a monolingual French model with the same performance would require about 82% of the non-embedding parameters. This hints there must be some effective capacity sharing between English and French.
Tokenizer fertility. According to Rust et al. (2021), fertility measures the average number of sub-words (tokens) produced per tokenized word. De facto, fertility has a theoretical lower bound of 1 which would imply that the tokenizer’s vocabulary contains every word in the corpus.
Appendix F Post-Mortem & Lessons learned
In line with our efforts of full transparency, we share some key lessons learned through this project, in light of a few months of hindsight and of recent developments posterior to the model release.
Tokenizer. In this work, we have developped our own tokenizer, inspired by Llama2 (Touvron et al., 2023b) but fitted to a bilingual French-English corpus to improve fertility on both languages. Several key decisions can impact tokenizer design. The vocabulary size will have large impacts on the parameter count of the model, potentially making embedding parameters a large share of the total parameter count especially for smaller models (Team et al., 2024a). One should experiment with the performance and memory trade-offs of using larger vocabulary sizes. Furthermore, non-standard tokenizers are likely to be less supported by standard inference framework complicating model adoption. Finally, we believe hand-designing a subset of tokens that are known to be useful at inference, for particular tasks, or have semantic coherence (common numbers, punctuation patterns, tokens corresponding to multiple choice templates or code syntax) probably degrades fertility but is useful in the long run from a performance and usability perspective.
Data Mix. It was always understood the data quality and quantity of the pretraining mix had large impacts on pretraining. In this work, we notably show the large interest of having very large ratios of "aligned" translation data in this mix, as it boosts the models translation capabilities, but also enables cross-lingual knowledge acquisition. Since CroissantLLM’s release, it has become understood that certain types of data heavily boost model reasoning capalities or performances on benchmarks. Typically, Abdin et al. (2024); Team et al. (2024b); Yang et al. (2024); Dubey et al. (2024), put a particular emphasis in including math and reasoning heavy corpora in the training corpus. Groeneveld et al. (2024) reports 24 point gains on MMLU scores by including more knowledge and reasoning rich sources to their training set, as well as improving data quality filtering. In order to source such data with highly "educative" values several approaches are possible, including distilling large model knowledge into fast text classifiers to identify high-quality content from Internet scale data (Penedo et al., 2023), or even generating diverse and high quality synthetic text using LLMs (Abdin et al., 2024; Ben Allal et al., 2024; Dubey et al., 2024). More generally, better pretraining and annealing data leads to better models, and many recent datasets would probably lead to better performance nowadays than the datamix we had constituted. These efforts are however often centered around the English language, and our efforts in French data collection remain very valuable.
Scheduler. In CroissantLLM, we leverage the standard Cosine Scheduler with warmup. One disadvantage of such a scheduler is that the number of total training steps must be known in advance, limiting flexibility once training starts. Inspired by the Vision literature (Zhai et al., 2022), the MiniCPM model (Hu et al., 2024) and several models since then have later confirmed the interest of infinite or WSD (Warmup-Stable-Decay) schedulers(Hägele et al., 2024). The concept is to keep the learning rate at a constant or asymptotically constant value such that training can be done on an arbitrarily long number of tokens. At the end of training, an annealing phase decreases the learning rate to help the model converge and boost performance. MiniCPM has further uncovered that annealing on a higher quality data mix is a very efficient strategy. Typically, by including knowledge-rich or reasoning-heavy data in this final training phase, large benchmark improvements are obtained. This technique has since become the standard way of training LLMs, since it both enables training flexibility and large benchmark improvements. We believe running an annealing phase on the CroissantLLM model would have largely boosted performance, especially if done with carefully selected high-quality data in French and English.
Leveraging Larger LLMs. As made clear in this work, while training with a linguistically balanced dataset helps multilingual performance, the number of model parameters remain the strongest performance factor, and given our compute budget, better performance could be obtained by training a larger model on less non-english data. This is perfectly in line with the findings of Kaplan et al. (2020a) with Chinchilla scaling laws. Models trained to be small and inference efficient such as CroissantLLM could however benefit from larger multilingual LLMs in multiple ways. Beyond the evergrowing use of LLM-generated data during pretraining (Abdin et al., 2024), small models can also benefit from larger LLMs as critics for alignement processes (Yang et al., 2024), but also to replace stochastic weight initialization by starting off from pruned-versions of larger models (Dubey et al., 2024). Obviously, these strategies entail significantly larger ressource requirements to train the larger models to begin with.
Model Release. Training a model is one thing, enabling people to use it is another. Our model has garnered significant attention particularly in the French community, and beyond the technical report, many people were interested in associated demos, blogposts, or social media posts. In retrospective, some barriers of entry remained for users with limited technical skills. The lack of day 0 support from third party inference libraries (Ollama [^72] https://ollama.com/, llama.cpp [^73] https://github.com/ggerganov/llama.cpp, MLCChat [^74] https://llm.mlc.ai/)has led to some community-proposed implementations that were suboptimal, leading to performance hits. Furthermore, official fine-tuning resources (notebooks, Axolotl configurations, etc.) could have largely helped users adapt CroissantLLM for specific use cases and drive adoption. Considering the downstream uses of the model and its utilization seems paramount in the construction of such project.
Building a model is an iterative process. Building upon the CroissantLLM work and the above listed insights, our team has since been able to train stronger models (Martins et al., 2024; Boizard et al., 2025). The field evolves quickly and getting everything right the first time is not easy. To build very good models, it seems important to us to construct the foundations iteratively, getting things to work at a small scale first, collecting some practical real-world feedback and keeping a core model development team stable.